OpenStudy (anonymous):

How would you solve sin^2x=sinx without graphing it?

OpenStudy (anonymous):

\[\sin^2(x) = (\sin(x))^2\] so dividing both sides by sin(x) gives sin(x) = 1

OpenStudy (anonymous):

It's just sin^2x=sinx, not =sinx^2

OpenStudy (anonymous):

Oh, NVM I see what ou meant.

OpenStudy (anonymous):

So then you would take the arcsin of 1?

OpenStudy (anonymous):

yes. gotta love brackets <3 :P

OpenStudy (anonymous):

lol

OpenStudy (asnaseer):

if you just divide by sin(x) then you will miss one of the solutions. you need to take the sin(x) from the right-hand-side to the left and factor as follows:\[\sin^2(x)=\sin(x)\]\[\therefore \sin^2(x)-\sin(x)=0\]\[\therefore \sin(x)(\sin(x)-1)=0\]solve from there

OpenStudy (anonymous):

Okay. How would you know if you are missing a sollution?

OpenStudy (asnaseer):

because you divided by sin(x) - that means you have thrown away one solution which is sin(x)=0

OpenStudy (anonymous):

Can you give me an example?

OpenStudy (asnaseer):

lets say you were given:\[x^2=x\]and you just divided both sides by x to get:\[x=1\]here you have missed the solution x=0

OpenStudy (asnaseer):

where as:\[x^2=x\]\[x^2-x=0\]\[x(x-1)=0\]\[x=0\qquad\text{or}\qquad x=1\]

OpenStudy (anonymous):

Okay, so you never divide by both sides when it comes to these types of problems, right?

OpenStudy (asnaseer):

correct

OpenStudy (anonymous):

Set it =0?

OpenStudy (anonymous):

Then simplify?

OpenStudy (anonymous):

asnaseer is right. my bad

OpenStudy (anonymous):

np.

OpenStudy (asnaseer):

no worries @Euler271 - we all make mistakes sometimes - that is what makes us human. :)

OpenStudy (anonymous):

Guys, can you go through with me on the next problem, tell me if I made a mistake?

OpenStudy (anonymous):

ya

OpenStudy (asnaseer):

sure - just post it as a new question please.

OpenStudy (anonymous):

Okay.